OpenStudy (anonymous):

Hi :) I'm taking a math course online and I can't understand this question. Please explain your steps, Thank You :D If f(x) = 2x^2 + 2, find f(5). a) 102 b)52 c)22 d)-1.5

OpenStudy (raden):

put x=5 to f(x), it means f(5) = 2(5)^2 + 2 f(5) = 2(25) + 2 f(5) = 50 + 2 f(5) = 52

OpenStudy (anonymous):

I got it:) Thank You!

OpenStudy (raden):

you're welcome